OpenStudy (anonymous):

solve: 5/(p-1)+1/(p^2-1)

OpenStudy (anonymous):

\[5/(p-1)+1/(p^2-1)\] \[=(5(p+1)+1)/(p^2-1)\] \[=(5p+6)/(p^2-1)\]